In reviewing activity for July, the controller of Mathis, Inc., collected the following data concerning direct materials.

 


Actual production
103,600units
Direct materials purchased (actual)$1,644,000
Standard cost of materials purchased
1,557,000
Standard direct materials costs per unit produced
14.20
Standard price times actual amount of materials used
1,408,950

 

Required:

Compute the direct materials cost variances. (Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

price variance 1,557,000- 1,644,000= 87,000 U


14.20* 103,600= 1,471,120

efficiency variance 1,471,120- 1,408,950= 62, 170 F


total DM cost variance 87,000- 62,170= 24,830 U